OpenStudy (anonymous):

A right triangle has one side with length 15 inches. It's hypotenuse measures 17 inches. what is the length of the third side?

OpenStudy (wolf1728):

Use Pythagorean Theorem hyp² = side 1² + side 2² side1² = hyp² -side2² side1² = 17² - 15² side1² = 289 -225 side1² = 64 side1 = 8

OpenStudy (anonymous):